# Eisenstein degeneration of Euler systems

### v1

@@ Abstract.
The aim of this note is to explore the Euler system of Beilinson–Kato elements in families passing through the critical $p$-stabilization of an Eisenstein series attached to two Dirichlet characters $(\psi,\tau)$. In this context, we establish an explicit connection with the system of circular units, utilizing suitable factorization formulas in a situation where several of the $p$-adic $L$-functions vanish. In that regard, our main results may be seen as an Euler system incarnation of the factorization formula of Bellaïche and Dasgupta. One of the most significant aspects is that, depending on the parity of $\psi$ and $\tau$, different phenomena arise; while some can be addressed with our methods, others pose new questions. Finally, we discuss analogous results in the framework of Beilinson–Flach classes.
@@ 1. Introduction
However, these local properties cannot be used when trying to compare the Beilinson–Kato system with circular units, as in this case no local conditions are involved at $p$. Therefore, this work should be seen as a natural continuation and an extension of the methods developed in [LRS24], where we modify our previous approach to cover the Beilinson–Kato case.
@@ 1.1. Set-up
Let $E_{r+2}(\psi,\tau)\in M_{k}(N,\xi)$ be an Eisenstein series of even weight $k=r+2$, level $N$ and nebentype $\xi=\psi\tau$, where $\psi$ and $\tau$ are primitive Dirichlet characters with conductors $N_{1}$ and $N_{2}$, respectively, with $N=N_{1}N_{2}$ and $\xi(-1)=(-1)^{r}$. Let ${\mathbf{f}}$ stand for the unique Coleman family passing through the critical slope $p$-stabilization of $f:=E_{r+2}(\psi,\tau)$, $f_{\beta}:=E_{r+2}^{\operatorname{crit}}(\psi,\tau)$. Unless explicitly stated otherwise, we will assume in the first part of the introduction that $r$ is even and that both $\tau$ and $\psi$ are even.
@@
The Coleman–Mazur eigencurve is smooth and locally étale at the point corresponding to $E_{r+2}^{\operatorname{crit}}(\psi,\tau)$ in the family ${\mathbf{f}}$.
This is known to hold in a high level of generality (see e.g. [LRS24, §6]), so this must be seen as a very mild assumption. From now on, let $X$ stand for a local uniformizer at that point. Similarly, we write $V({\mathbf{f}})^{*}$ (resp. $V^{c}({\mathbf{f}})^{*}$) for the non-compactly supported (resp. compactly supported) cohomology of ${\mathbf{f}}$. Let $\mathcal{H}_{\Gamma}$ be the distribution algebra of the cyclotomic extension $\Gamma\cong\mathbf{Z}_{p}^{\times}$, and ${\mathbf{j}}\colon\Gamma\hookrightarrow\mathcal{H}_{\Gamma}^{\times}$ the universal character, which we regard as a character of the Galois group by composition with the cyclotomic character.
To formulate the main result of this note, we need to introduce the following two Euler systems:
@@
    The Kato Euler system attached to a Coleman family, which is a class
@@
Letting $\Gamma^{+}=\Gamma/\langle\pm 1\rangle$, and ${\mathbf{j}}_{+}$ for its universal character, we can see the class in the smaller subspace $H^{1}(\mathbf{Q},\mathbf{Q}_{p}(\tau^{-1})(1+r)\otimes\mathcal{H}_{\Gamma}(-{\mathbf{j}}_{+}))$, since the Iwasawa cohomology of the other half of the weight space vanishes. Our comparison of Beilinson–Kato classes and circular units will involve three different kinds of $p$-adic $L$-functions:
@@
    The (one-variable) Kubota–Leopoldt $p$-adic $L$-function $L_{p}(\chi)$ attached to a Dirichlet character $\chi$.
@@ 1.2. Main results
The first main result of this article may be seen as a direct comparison between the Kato Euler system and that of circular units. In order to state it as simple as possible in the introduction, we will assume that the leading term in the argument of Proposition 5.12 corresponds to $n=0$. This means that we assume that $\kappa(f_{\beta})$ is equal to the class $\kappa^{*}(f_{\beta})$ introduced later in the text. For a more precise version of the result, see 5. For the moment, we need an extra assumption.
@@
The adjoint $p$-adic $L$-function $L_{p}(\operatorname{Ad}{\mathbf{f}})$ has a simple zero at the critical Eisenstein point corresponding to $E_{r+2}^{\operatorname{crit}}(\psi,\tau)$.
It is known, by the results of Bellaïche, [Bel21, §9], that $L_{p}(\operatorname{Ad}{\mathbf{f}})$ vanishes at $f_{\beta}$, but we do not know whether or not it presents a simple zero. The result we state also involves the logarithmic distribution: for a continuous character $\sigma:\mathbf{Z}_{p}^{\times}\rightarrow\mathbf{C}_{p}$, the function $\frac{d^{k}\sigma}{dz^{k}}\cdot\frac{z^{k}}{\sigma(z)}$ is a constant, and $\log^{[k]}\in\mathbf{C}_{p}$ is defined to be this constant. Further, we write $\mathcal{E}_{N}$ for the two-variable function corresponding to the product of bad Euler factors introduced in Definition 4.7. Finally, let $\lambda$ denote the function defined over the Iwasawa algebra such that its specialization at a character $\sigma$ corresponding to an integer $s$ and a Dirichlet character $\chi$ of $p$-power conductor is $\chi^{-1}(N_{2})\langle N_{2}\rangle^{-s+1}$.
@@
Suppose that $r$ is even and that the characters $\psi$ and $\tau$ are both even. Then, under Assumption 1.1 and Assumption 1.2, the following equality holds in $H^{1}(\mathbf{Q},\mathbf{Q}_{p}(\tau^{-1})(1+r)\otimes\mathcal{H}_{\Gamma}(-{\mathbf{j}})))$:
$\kappa(f_{\beta})=C\cdot\lambda\cdot\mathcal{E}_{N}\cdot\frac{\log^{[r+1]}\cdot L_{p}^{\prime}(f_{\beta},\chi)\cdot L_{p}(\psi,{\mathbf{j}}+1)}{L_{p}^{\prime}(\operatorname{Ad}f_{\beta})}c(\tau)(r),$
where $C$ is a non-zero constant which does not depend on $j$ and $L_{p}^{\prime}(f_{\beta},\chi)$ stands for the derivative of the two-variable $p$-adic $L$-function $L_{p}({\mathbf{f}},\chi)$ restricted to the line $s=k/2$, and evaluated at $f_{\beta}$.
The constant $C$ arises during the interpolation of Ohta’s canonical differentials, and it may be seen as the less understood piece of the picture; roughly speaking, it has to do with the interpolation of the Eichler–Shimura isomorphisms around critical Eisenstein points. Observe that the factors $L_{p}^{\prime}(f_{\beta},\chi)$ and $L_{p}^{\prime}(\operatorname{Ad}f_{\beta})$ only depend on the weight, but we have opted to include them separately to emphasize this dependence. This means that we have the equality
$\kappa(f_{\beta})=\tilde{C}\cdot\lambda\cdot\mathcal{E}_{N}\cdot\log^{[r+1]}\cdot L_{p}(\psi,{\mathbf{j}}+1)\cdot c(\tau)(r),$
where $\tilde{C}$ is an arbitrary constant that can potentially be zero.
@@
    Deformation of the Beilinson–Kato elements in families passing through a critical Eisenstein point. The main result in this direction is stated as Theorem 5.1.
@@
    Interpolation of the canonical differentials $\eta_{{\mathbf{f}}}$ and $\omega_{{\mathbf{f}}}$ in families, together with a development of the explicit reciprocity laws in the critical Eisenstein situation. The former is done in Section 2.3, and the latter in Theorem 5.8.
@@
    Bellaïche–Dasgupta’s factorization formula [BD15], expressing the $p$-adic $L$-function of a critical Eisenstein series as the product of two Kubota–Leopoldt $p$-adic $L$-functions.
The odd case. If $r$ is even, $\tau$ and $\psi$ are both odd, and we further assume that $\tau(p)\neq 1$, one can directly claim that $\kappa(f_{\beta})=0$, since Iwasawa cohomology is zero along that half of the weight space, i.e., $H^{1}(\mathbf{Q},\mathbf{Q}_{p}(\tau^{-1})(1+r)\otimes\mathcal{H}_{\Gamma}(-{\mathbf{j}}_{+}))$. This means that the class $\kappa({\mathbf{f}})$ lifts to an element in $H^{1}(\mathbf{Q},V({\mathbf{f}})^{*}(-{\mathbf{j}}_{+}))$. In this case, we can project $V({\mathbf{f}})^{*}$ onto the cohomology of the quotient
@@
This class belongs to the one-dimensional space spanned by the circular unit $c(\psi)(-1)$. Unfortunately, our methods are not enough to get a good description of that class.
@@
$L_{p}^{\prime}(f_{\beta})(s)=\lambda\cdot N_{2}^{-s}\log^{[r+1]}L_{p}(\psi,s)\cdot L_{p}(\tau,s-r-1),$
@@
The organization of this note is as follows: we begin with an introduction of the Galois representations and periods involved in our comparison, with emphasis on the interpolation of the canonical differentials in a neighborhood of a critical Eisenstein point. Next, we fix our conventions for the systems of circular units and Kato elements, discussing the vanishing of the $p$-adic $L$-function of the adjoint and one of the Kitagawa–Mazur $p$-adic $L$-functions. This allows us to develop the study of the degeneration of Beilinson–Kato elements and carry out the comparison between both Euler systems. Section 6 serves to discuss the limitation of our methods in the cases where the characters $\psi$ and $\tau$ are not both odd, and raises some interesting questions that we leave for forthcoming work; more specifically, there are two main unsolved issues: (a) the description of $\kappa(f_{\beta})_{2}$; (b) the extension of the factorization formula of [BD15]. We end up the note with an analogy with the simpler case of Beilinson–Flach elements.
@@ 1.3. Relation with other works
A similar idea in the framework of congruences among modular forms was explored in [RR21]. In that situation, the modulo $p$ congruences suggested the possibility of an extra vanishing in certain cases, which led us to analyze a second derivative with techniques that went beyond the explicit reciprocity laws. It would be interesting to study if those results, mainly due to Sharifi and Fukaya–Kato, admit a counterpart in this scenario.
The idea of using degeneration of Eisenstein series is not new in the theory of modular forms. For instance, in the setting of irregular weight one modular forms, when one takes the Eisenstein series of characters $(\chi_{1},\chi_{2})$, with $\chi_{1}(p)=\chi_{2}(p)$, there are three (ordinary) families passing through it: the families of Eisenstein series ${\mathbf{E}}(\chi_{1},\chi_{2})$ and ${\mathbf{E}}(\chi_{2},\chi_{1})$, but also a third family which is generically cuspidal. This idea is explored in the work of [BDV22] to prove Perrin-Riou’s conjecture via a comparison between Beilinson–Kato systems and Heegner cycles.
@@
Finally, we expect to extend the study of these phenomena to settings beyond $\operatorname{GL}_{2}$, like the Siegel case, where one can study endoscopy phenomena emerging in the theory of Saito–Kurokawa and Yoshida lifts.
@@ 1.4. Acknowledgements
The authors would like to thank Kazim Büyükboduk, Samit Dasgupta, David Loeffler, Alexandre Maksoud, Alice Pozzi, and Victor Rotger for informative conversations around the topic of this article. The second author is especially grateful to David Loeffler, with whom he learned many of the techniques developed in this paper during the writing of the prequel [LRS24] and the subsequent discussions around the topic.
@@ 2.1. Eisenstein series and Galois representations
Let $p>2$ be a fixed prime number. Let $\psi,\tau$ be Dirichlet characters and $r\geqslant 0$ an integer with $\psi\tau(-1)=(-1)^{r}$. Let $N_{1}$ and $N_{2}$ denote the conductors of $\psi$ and $\tau$, respectively, and let $N=N_{1}N_{2}$. Choose an embedding $\mathbf{Q}(\psi,\tau)\hookrightarrow L$, where $\mathbf{Q}(\psi,\tau)\subset\mathbf{C}$ is the finite extension of $\mathbf{Q}$ generated by the values of $\psi$ and $\tau$, and $L$ is a finite extension of $\mathbf{Q}_{p}$.
We write $f=E_{r+2}(\psi,\tau)$ for the classical modular form of level $N$ and weight $r+2$ given by
@@
where ()(*)( ∗ ) is the constant defined e.g in [DS05, Ch. 4]. In particular, it is equal to $\delta(\psi)L(\tau,1-k)/2$, where $\delta(\psi)=1$ if $\psi$ is trivial and 0 if not, and $L(\tau,\cdot)$ is the Dirichlet $L$-series attached to the character $\tau$. Note that we may regard the $q$-expansion coefficients $a_{n}(f)$ of $f=E_{r+2}(\psi,\tau)$ as elements of $L$.
@@
and similarly for $f_{\alpha}(q)$. We say $f_{\beta}$ is non-critical if the generalised eigenspace of overconvergent modular forms $M_{r+2}^{{\dagger}}(\Gamma)_{(T=f_{\beta})}$ associated to $f_{\beta}$ is 1-dimensional.
The next definition plays a crucial role for establishing the properties of the Galois representations.
@@
  • Let $\ell$ be a prime. We write $\operatorname{Frob}_{\ell}$ for an arithmetic Frobenius element at $\ell$ in $\operatorname{Gal}(\overline{\mathbf{Q}}/\mathbf{Q})$; this depends on the choice of an embedding of $\overline{\mathbf{Q}}$ into $\overline{\mathbf{Q}}_{\ell}$, and is well-defined modulo the inertia subgroup for this embedding.

### v3

# Eisenstein degeneration of Beilinson–Kato classes and circular units

@@ Abstract.
The aim of this note is to explore the Euler system of Beilinson–Kato elements in families passing through the critical $p$-stabilization of an Eisenstein series attached to two Dirichlet characters $(\psi,\tau)$, a setting which is not covered by the recent works of Benois and Büyükboduk. In this context, we establish an explicit connection with the system of circular units, utilizing suitable factorization formulas in a situation where several of the $p$-adic $L$-functions vanish. In that regard, our main results may be seen as an Euler system incarnation of the factorization formula of Bellaïche and Dasgupta. One of the most significant aspects is that, depending on the parity of $\psi$ and $\tau$, different phenomena arise; while some can be addressed with our methods, others pose new questions. Finally, we discuss analogous results in the framework of Beilinson–Flach classes.
@@ 1. Introduction
However, these local properties cannot be used when trying to compare the Beilinson–Kato system with circular units, as in this case no local conditions are involved at $p$. Therefore, this work should be seen as a natural continuation and an extension of the methods developed in [LRS24], where we modify our previous approach to cover the Beilinson–Kato case. This case turns out to be ostensibly different, showing in an easier instance some of the properties already envisaged in [BB24] about the behavior of this Euler system. Further, the main result can be interpreted as an Euler-system incarnation of the results of Bellaïche–Dasgupta.
@@ 1.1. Set-up
Let $E_{r+2}(\psi,\tau)\in M_{k}(N,\xi)$ be an Eisenstein series of even weight $k=r+2$, level $\Gamma_{1}(N)$ and nebentype $\xi=\psi\tau$, where $\psi$ and $\tau$ are primitive Dirichlet characters with conductors $N_{1}$ and $N_{2}$, respectively, with $N=N_{1}N_{2}$ and $\xi(-1)=(-1)^{r}$. Let $\mathbf{f}$} stand for the unique Coleman family passing through the critical slope $p$-stabilization of $f:=E_{r+2}(\psi,\tau)$, $f_{\beta}:=E_{r+2}^{\operatorname{crit}}(\psi,\tau)$. Unless explicitly stated otherwise, we will assume in the first part of the introduction that $r$ is even and that both $\tau$ and $\psi$ are even.
@@
The critical slope $p$-stabilization $f_{\beta}$ is non-crticial, i.e., the generalised eigenspace of overconvergent modular forms associated to $f_{\beta}$ is 1-dimensional. Further, the Coleman–Mazur eigencurve is smooth and locally étale at the point corresponding to $E_{r+2}^{\operatorname{crit}}(\psi,\tau)$ in the family $\mathbf{f}$}.
This is known to hold in a high level of generality (see e.g. [LRS24, §A6]), so this must be seen as a very mild assumption. We later discuss several equivalent formulations for this condition. From now on, let $X$ stand for a local uniformizer at that point. Similarly, we write $V({\mathbf{f}})^{*}$ (resp. $V^{c}({\mathbf{f}})^{*}$) for the non-compactly supported (resp. compactly supported) cohomology of $\mathbf{f}$}. Let $\mathcal{H}_{\Gamma}$ be the distribution algebra of the cyclotomic extension $\Gamma\cong\mathbf{Z}_{p}^{\times}$, i.e., the global sections of the rigid analytic space associated with $\mathrm{Spf}\,\mathbf{Z}_{p}[\![\Gamma]\!]$. Let ${\mathbf{j}}\colon\Gamma\hookrightarrow\mathcal{H}_{\Gamma}^{\times}$ the universal character, which we regard as a character of the Galois group by composition with the cyclotomic character.
To formulate the main result of this paper, we need to introduce the following two Euler systems.
@@
    The Kato Euler system (sometimes called Beilinson–Kato Euler system) attached to a Coleman family, which is a class
@@
Letting $\Gamma^{+}=\Gamma/\langle\pm 1\rangle$, and writing $\mathbf{j}$$_{+}$ for its universal character, we can see the class in the smaller subspace $H^{1}(\mathbf{Q},\mathbf{Q}_{p}(\tau^{-1})(1+r)\otimes\mathcal{H}_{\Gamma}(-{\mathbf{j}}_{+}))$, since the Iwasawa cohomology of the other half of the weight space vanishes. Our comparison of Beilinson–Kato classes and circular units will involve three different kinds of $p$-adic $L$-functions:
@@
    The (one-variable) Kubota–Leopoldt $p$-adic $L$-function $L_{p}(\chi)$ attached to an even and non-trivial Dirichlet character $\chi$.
@@ 1.2. Main results
The first main result of this article may be seen as a direct comparison between the Kato Euler system and that of circular units. In order to state it as simple as possible in the introduction, we will assume that the leading term in the argument of Proposition 5.13 corresponds to $n=0$. This means that we assume that $\kappa(f_{\beta})$ is equal to the class $\kappa^{*}(f_{\beta})$ introduced later in the text. For a more precise version of the result, see 5. The Kato Euler system depends on the choice of an odd Dirichlet character $\chi$, and under the current sign assumptions, $L_{p}({\mathbf{f}},\chi)$ vanishes at $(f_{\beta},r/2+1)$. Write $L_{p}^{\prime}(f_{\beta},\chi)$ for the derivative of the two-variable $p$-adic $L$-function $L_{p}({\mathbf{f}},\chi)$ restricted to the line $s=k/2$, and evaluated at $f_{\beta}$.
For the moment, we need an extra set of assumptions, which assert that the derivatives of the relevant $p$-adic $L$-functions do not vanish. The second part of the assumption may be relaxed, as we discuss in the body of the text.
@@
The adjoint $p$-adic $L$-function $L_{p}(\operatorname{Ad}{\mathbf{f}})$ has a simple zero at the critical Eisenstein point corresponding to $E_{r+2}^{\operatorname{crit}}(\psi,\tau)$. Moreover, $L_{p}^{\prime}(f_{\beta},\chi)$ is non-zero.
It is known, by the results of Bellaïche, [Bel21, §9], that $L_{p}(\operatorname{Ad}{\mathbf{f}})$ vanishes at $f_{\beta}$, but we do not know whether or not it presents a simple zero. The result we state also involves the logarithmic distribution: for a continuous character $\sigma:\mathbf{Z}_{p}^{\times}\rightarrow\mathbf{C}_{p}$, the function $\frac{d^{k}\sigma}{dz^{k}}\cdot\frac{z^{k}}{\sigma(z)}$ is a constant, and $\log^{[k]}\in\mathbf{C}_{p}$ is defined to be this constant. In addition, we write $\mathcal{E}_{N}$ for the two-variable function corresponding to the product of the bad Euler factors introduced in Definition 4.8. Finally, let $\lambda$ denote the function defined over the Iwasawa algebra such that its specialization at a character $\sigma\colon t\mapsto t^{s}\xi(s)$ corresponding to an integer $s$ and a Dirichlet character $\xi$ of $p$-power conductor is $\xi^{-1}(N_{2})\langle N_{2}\rangle^{-s+1}$.
@@
Suppose that $r$ is even and that the characters $\psi$ and $\tau$ are both even. Then, under Assumption 1.1 and Assumption 1.2, the following equality holds in $H^{1}(\mathbf{Q},\mathbf{Q}_{p}(\tau^{-1})(1+r)\otimes\mathcal{H}_{\Gamma}(-{\mathbf{j}}))$:
$\kappa(f_{\beta})=\left(\frac{C\cdot\lambda\cdot\mathcal{E}_{N}\cdot L_{p}^{\prime}(f_{\beta},\chi)}{L_{p}^{\prime}(\operatorname{Ad}f_{\beta})}\right)\cdot L_{p}(\psi,{\mathbf{j}}+1)\cdot c(\tau)(r),$
where $C$ is a constant which does not depend on $j$.
The constant $C$ arises during the interpolation of Ohta’s canonical differentials, and it may be seen as the less understood piece of the picture; roughly speaking, it has to do with the interpolation of the Eichler–Shimura isomorphisms around critical Eisenstein points. Observe that the factors $L_{p}^{\prime}(f_{\beta},\chi)$ and $L_{p}^{\prime}(\operatorname{Ad}f_{\beta})$ only depend upon the weight, but we have opted to include them separately to emphasize this dependence. This means that we have the equality
$\kappa(f_{\beta})=\tilde{C}\cdot\lambda\cdot\mathcal{E}_{N}\cdot L_{p}(\psi,{\mathbf{j}}+1)\cdot c(\tau)(r),$
where $\tilde{C}$ is an arbitrary constant. In particular, it can be zero.
@@
    Deformation of the Beilinson–Kato elements in families passing through a critical Eisenstein point. The main result in this direction is stated as Theorem 5.1. Observe that the critical Eisenstein situation requires a careful study of the variation in Coleman families, in a setting where the arguments of [BB22] and [BB24] do not directly work.
@@
    Interpolation of the canonical differentials $\eta_{{\mathbf{f}}}$ and $\omega_{{\mathbf{f}}}$ in families, together with a development of the explicit reciprocity laws around the critical Eisenstein point. The former is done in Section 2.3, and the latter in Theorem 5.9.
@@
    Bellaïche–Dasgupta’s factorization formula [BD15], expressing the $p$-adic $L$-function of a critical Eisenstein series as the product of two Kubota–Leopoldt $p$-adic $L$-functions and a logarithmic term.
The odd case. If $r$ is even, $\tau$ and $\psi$ are both odd, and we further assume that $\tau(p)\neq 1$, one can directly claim that $\kappa(f_{\beta})=0$, since Iwasawa cohomology is zero along that half of the weight space, i.e., $H^{1}(\mathbf{Q},\mathbf{Q}_{p}(\tau^{-1})(1+r)\otimes\mathcal{H}_{\Gamma}(-{\mathbf{j}}_{+}))=0$. This means that the class $\kappa({\mathbf{f}})$ lifts to an element in $H^{1}(\mathbf{Q},V({\mathbf{f}})^{*}(-{\mathbf{j}}_{+}))$. In this case, we can project $V({\mathbf{f}})^{*}$ onto the cohomology of the quotient
@@
This class belongs to the one-dimensional space spanned by the circular unit Euler system $c(\psi)(-1)$. Unfortunately, our methods are not enough to get a good description of that class, and this will be developed in the sequel to this paper, extending the theory we have developed so far. This may be seen as an instance of exceptional zero phenomenon, where the vanishing, however, is not caused by a zero in an Euler factor.
@@
$L_{p}^{\prime}(f_{\beta})(s)=\lambda N_{2}^{-s}\log^{[r+1]}\cdot L_{p}(\psi,s)\cdot L_{p}(\tau,s-r-1),$
@@
The organization of this note is as follows: we begin with an introduction of the Galois representations and periods involved in our comparison, with emphasis on the interpolation of the canonical differentials in a neighborhood of a critical Eisenstein point. Next, we fix our conventions for the systems of circular units and Kato elements, discussing the vanishing of the $p$-adic $L$-function of the adjoint and one of the Kitagawa–Mazur $p$-adic $L$-functions. This allows us to develop the study of the degeneration of Beilinson–Kato elements and carry out the comparison between both Euler systems. Section 6 serves to discuss the limitation of our methods in the cases where the characters $\psi$ and $\tau$ are not both odd, and raises some interesting questions that we leave for forthcoming work; more specifically, there are two main unsolved issues: (a) the description of $\kappa(f_{\beta})_{2}$; (b) the extension of the factorization formula of [BD15]. We end up the note with an analogy with the simpler case of Beilinson–Flach elements, which is further developed in a sequel of this note, where we extend the methods of this article.
In Appendix A, we discuss in detail the construction of rank-$2$ big Galois representations. The materials presented in this appendix should be well known to the experts, but we were unable to find a detailed reference. We also note that the appendix is written in a way that suggests a possible generalization to higher-rank groups.
@@ 1.3. Relation with other works
A similar idea in the framework of congruences among modular forms was explored in [RR21]. In that situation, the modulo $p$ congruences suggested the possibility of an extra vanishing in certain cases, which led us to analyze a second derivative with techniques that went beyond the explicit reciprocity laws. It would be interesting to study whether those results, mainly due to Sharifi and Fukaya–Kato, admit a counterpart in this scenario.
The idea of using the degeneration of Eisenstein series is not new in the theory of modular forms. For instance, in the setting of irregular weight one modular forms, when one takes the Eisenstein series of characters $(\chi_{1},\chi_{2})$, with $\chi_{1}(p)=\chi_{2}(p)$, there are three (ordinary) families passing through it: the families of Eisenstein series ${\mathbf{E}}(\chi_{1},\chi_{2})$ and ${\mathbf{E}}(\chi_{2},\chi_{1})$, but also a third family which is generically cuspidal. This idea is explored in the work of [BDV22] to prove Perrin-Riou’s conjecture via a comparison between Beilinson–Kato systems and Heegner cycles.
@@
Finally, we expect to extend the study of these phenomena to settings beyond $\operatorname{GL}_{2}$, such as the Siegel case, where one can study endoscopy phenomena emerging in the theory of Saito–Kurokawa and Yoshida lifts.
@@ 1.4. Acknowledgements
We would like to thank Kazim Büyükboduk for very helpful conversations and for his comments on an earlier version of the manuscript. The second author is especially grateful to David Loeffler, with whom he learned many of the techniques developed in this paper during the writing of the prequel [LRS24] and the subsequent discussions around the topic. The authors extend their gratitude to Samit Dasgupta, Alexandre Maksoud, Alice Pozzi, and Victor Rotger for informative conversations and correspondence around the topic of this article. J.P. and O.R. express their deepest gratitude to Ju-Feng Wu for writing the appendix to this paper.
@@ 2.1. Eisenstein series and Galois representations
Let $p>2$ be a fixed prime number. Let $\psi,\tau$ be Dirichlet characters and $r\geqslant 0$ an integer with $\psi\tau(-1)=(-1)^{r}$. Let $N_{1}$ and $N_{2}$ denote the conductors of $\psi$ and $\tau$, respectively, and let $N=N_{1}N_{2}$. Throughout this article, we assume that $p\nmid N$. Choose an embedding $\mathbf{Q}(\psi,\tau)\hookrightarrow L$, where $\mathbf{Q}(\psi,\tau)\subset\mathbf{C}$ is the finite extension of $\mathbf{Q}$ generated by the values of $\psi$ and $\tau$, and $L$ is a finite extension of $\mathbf{Q}_{p}$.
We write $f=E_{r+2}(\psi,\tau)$ for the classical modular form of level $\Gamma_{1}(N)$ and weight $r+2$ given by
@@
where ()(*) is the constant defined e.g in [DS05, Ch. 4]. In particular, it is equal to $\delta(\psi)L(\tau,1-k)/2$, where $\delta(\psi)=1$ if $\psi$ is trivial and 0 if not, and $L(\tau,\cdot)$ is the Dirichlet $L$-series attached to the character $\tau$. Note that we may regard the $q$-expansion coefficients $a_{n}(f)$ of $f=E_{r+2}(\psi,\tau)$ as elements of $L$.
@@
and similarly for $f_{\alpha}(q)$.
Definition 2.1.
We say $f_{\beta}$ is non-critical if the generalised eigenspace of overconvergent modular forms $M_{r+2}^{{\dagger}}(\Gamma)_{(T=f_{\beta})}$ associated to $f_{\beta}$ is 1-dimensional.
In [BD15, Rk. 1.5] the authors discuss how to reinterpret the criticality condition in terms of the non-vanishing of a $p$-adic $L$-value. The next definition plays a crucial role for establishing the properties of the Galois representations.
@@
  • Let $\ell$ be a prime. We write $\operatorname{Frob}_{\ell}$ for an arithmetic Frobenius element at $\ell$ in $\operatorname{Gal}(\overline{\mathbf{Q}}/\mathbf{Q})$; this depends on the choice of an embedding of $\overline{\mathbf{Q}}$ into $\overline{\mathbf{Q}}_{\ell}$, and is well defined modulo the inertia subgroup for this embedding.